### v2

# ELLA: Efficient Lifelong Learning for Adapters in Large Language Models

Abstract
Large Language Models (LLMs) suffer severe catastrophic forgetting when adapted sequentially to new tasks in a continual learning (CL) setting. Existing approaches are fundamentally limited: replay-based methods are impractical and privacy-violating, while strict orthogonality-based methods collapse under scale: each new task is projected onto an orthogonal complement, progressively reducing the residual degrees of freedom and eliminating forward transfer by forbidding overlap in shared representations. In this work, we introduce ELLA, a training framework built on the principle of selective subspace de-correlation. Rather than forbidding all overlap, ELLA explicitly characterizes the structure of past updates and penalizes alignments along their high-energy, task-specific directions, while preserving freedom in the low-energy residual subspaces to enable transfer. Formally, this is realized via a lightweight regularizer on a single aggregated update matrix. We prove this mechanism corresponds to an anisotropic shrinkage operator that bounds interference, yielding a penalty that is both memory- and compute-constant regardless of task sequence length. ELLA requires no data replay, no architectural expansion, and negligible storage. Empirically, it achieves state-of-the-art CL performance on three popular benchmarks, with relative accuracy gains of up to $9.6\%$ and a $35\times$ smaller memory footprint. Further, ELLA scales robustly across architectures and actively enhances the model’s zero-shot generalization performance on unseen tasks, establishing a principled and scalable solution for constructive lifelong LLM adaptation.
ELLA: Efficient Lifelong Learning for Adapters in Large Language Models
Shristi Das Biswasthanks: Work conducted at Amazon. Purdue University          Yue Zhang AWS          Anwesan Pal AWS AI Labs          Radhika Bhargava AWS          Kaushik Roy Purdue University
11footnotetext: https://sites.google.com/view/ella-llm/home.
1 Introduction
Large Language Models (LLMs) have demonstrated remarkable generalization capabilities across a wide range of downstream tasks, largely attributed to their large-scale pretraining on diverse corpora Brown et al. (2020); Touvron et al. (2023); Achiam et al. (2023). As LLMs are increasingly deployed in real-world applications, they must be continuously adapted to evolving user needs and task distributions for long-term practical deployment. This setting, commonly studied under the paradigm of continual learning (CL), requires models to acquire new knowledge sequentially without costly full-retraining Ruvolo and Eaton (2013). However, sequential finetuning of LLMs remains highly susceptible to catastrophic forgetting (CF) - the tendency to overwrite prior knowledge when new tasks are introduced McCloskey and Cohen (1989), and loss of plasticity - deterioration in the ability to learn new information over time Dohare et al. (2021). These challenges are especially pronounced in rehearsal-free settings, where previously seen data cannot be stored due to privacy or storage constraints Chaudhry et al. (2019).
A promising line of work leverages parameter-efficient fine-tuning (PEFT) strategies, such as Low-Rank Adaptation (LoRA) Hu et al. (2022), to reduce the overhead of task-specific adaptation. Adapter-based approaches restrict updates to a small set of trainable parameters while leaving the base model frozen, making them a natural fit for lower-compute CL Qin and Joty (2021); Song et al. (2023). Yet, even with such modularity, sequential adapter training often yields severe forgetting when past knowledge is not revisited Wang et al. (2024b). Prior efforts to address this include expanding model capacity Wang et al. (2024a), isolating weights to reduce interference Aljundi et al. (2017); Li et al. (2019); Wang et al. (2023c), subspace orthogonality Wang et al. (2023a); Liao et al. (2025), or gradient projection from previous tasks Qiao and Mahdavi (2024). While effective to varying degrees, these often either limit forward knowledge transfer from past tasks, add storage overhead, or overlook activation-level interference, where forgetting actually manifests Ke et al. (2021).
A truly effective continual learning framework should balance knowledge retention with representation reuse, preventing harmful interference while allowing useful subspaces to be shared for new tasks Wu et al. (2024). In practice, not all overlap is harmful – low-magnitude directions from past tasks may encode generic linguistic or semantic patterns that can accelerate learning on future tasks if safely reused. Unfortunately, existing CL methods either eliminate all overlap or fuse past knowledge through heavyweight mechanisms like controller networks or rank-conditioned fusion, which limits scalability and increases complexity Zhao et al. (2024); Liu et al. (2023); Liao et al. (2025).
In this work, we propose ELLA, a simple and scalable CL framework for LLMs that mitigates forgetting without relying on replay buffers, parameter storage overheads or additional routing heuristics. ELLA introduces a subspace-aware regularization strategy that operates directly in weight space: we track the representational subspaces induced by past adapters and penalize updates that cause the new task’s adapter to align too closely with them. This cross-task de-correlation simultaneously encourages task-specific specialization by preserving prior representational geometry to reduce interference, while permitting the reuse of low-magnitude directions to enable forward transfer. As we formally prove in Appendix A, this selective regularization corresponds to an anisotropic shrinkage operator that provably bounds interference, providing a theoretical foundation for ELLA’s ability to balance stability and plasticity.
Through extensive experiments on the Standard CL Benchmark Zhang et al. (2015), Long Sequence Benchmark Razdaibiedina et al. (2023) and TRACE Wang et al. (2023b), we demonstrate that ELLA achieves state-of-the-art performance while scaling effectively across model sizes (from $770M$ to $8B$). Furthermore, ELLA generalizes well across architecture families (e.g., T5 Raffel et al. (2020), LLaMA Touvron et al. (2023)), and unlike prior work, uniquely improves generalization to unseen tasks. ELLA does not rely on task identities during inference, and is hence naturally compatible with the instruction-tuning paradigm Wang et al. (2022a), thereby preserving the generalization capabilities of LLMs in zero-shot and open-ended settings. Notably, our method can also be seamlessly integrated with existing continual learning methods to further enhance their effectiveness, without requiring additional supervision or auxiliary components. Moreover, ELLA is architecture-agnostic, requires no access to past data, and introduces negligible computational or memory overhead. In summary, our contributions are as follows:
  • We propose ELLA, a replay-free plug-and-play CL framework for LLMs that balances plasticity and mitigates CF via subspace-aware regularization, and we provide a formal theoretical analysis of its properties.
  • We provide extensive empirical evidence that ELLA establishes a new state-of-the-art in both performance and efficiency across three popular CL benchmarks, decisively outperforming prior methods without task labels, replay, or additional overhead.
  • We show that ELLA scales effectively across architectures and maintains strong generalization to unseen tasks, highlighting its practical advantages for real-world deployment.
2 Related Works
Continual Learning   The primary challenge in CL is to balance model stability (resisting catastrophic forgetting) with plasticity (acquiring new knowledge). Prior approaches to address this problem can be broadly categorized by their core mechanism, each presenting a distinct set of trade-offs. (i) Rehearsal-based methods store a subset of past data and interleave it during subsequent training phases. These include strategies such as experience replay Riemer et al. (2018) or constrained optimization Aljundi et al. (2017); Chaudhry et al. (2019); He et al. (2024) that jointly train on current and previous samples. While empirically strong, this approach is often impractical due to significant storage overhead and potential violations of data privacy constraints. (ii) Regularization-based methods add penalty terms to the loss that restrict updates to parameters critical for old tasks Du et al. (2024); Li and Hoiem (2017); de Masson D’Autume et al. (2019). For instance, Kirkpatrick et al. (2017) slows updates on important weights, while Farajtabar et al. (2020) constrains new updates to be orthogonal to gradients of old tasks. These methods depend on importance metrics that are often brittle and prevent any forward transfer, limiting adaptability to new tasks. (iii) Architecture-based methods reduce interference through task-specific modules or by expanding model size Li et al. (2019); Wang et al. (2023c). While Razdaibiedina et al. (2023) appends a new learned soft prompt per task,  Qin and Joty (2021) utilizes a large soft prompt that is continuously trained on all tasks. While this effectively isolates task knowledge, it typically results in a linear growth in parameter count with the number of tasks, posing significant scalability challenges and often requiring explicit task labels at inference.
Parameter-Efficient Continual Learning   Parameter-Efficient Fine-Tuning (PEFT) has emerged as a highly promising substrate for CL, with Low-Rank Adaptation (LoRA) Hu et al. (2022) being a particularly effective technique. However, naïvely applying LoRA sequentially leads to significant forgetting owing to interference across tasks. As tasks arrive sequentially, the adapter weights $A_{t}$, $B_{t}$ learned for a new task $t$ are trained from scratch without awareness of previous tasks’ LoRA subspaces. Consequently, overlapping update directions in parameter space may inadvertently override past knowledge, especially under limited parameter budgets. Without coordination across tasks, these LoRA components can destructively interfere in spaces crucial for previous tasks, degrading performance on earlier tasks while optimizing for the current one. Recent works in CL literature addressing this issue primarily falls into two schools of thought.
The first approach applies strict orthogonality to enforce zero overlap between the LoRA subspaces of different tasks Wang et al. (2023a); Qiao and Mahdavi (2024). While this hard constraint effectively mitigates interference, it imposes an overly restrictive inductive bias. By precluding any form of representation sharing, it severely inhibits beneficial forward transfer and prevents reuse of low-importance shared space across related tasks. Worse still, as tasks accumulate, the available orthogonal space quickly diminishes, leading to progressively less capacity for future learning. This rigid decoupling causes inefficient use of adapter capacity, hindering the goal of cumulative learning. Furthermore, storing all past LoRA weights incurs high memory overhead that scales linearly with the number of tasks, further limiting scalability.
The second approach involves more complex fusion and replay mechanisms. Methods like DATA Liao et al. (2025) or Recurrent KIF Feng et al. (2025) often rely on replay buffers, multiple LoRA adapters per task, or intricate gating mechanisms to learn how to compose adapters. These methods reintroduce significant architectural complexity and computational overhead, sacrificing the very simplicity and efficiency that make PEFT an attractive paradigm for CL in the first place.
Differently, ELLA charts a third path. The prevailing paradigms are caught in a dilemma: rehearsal is impractical, naïve regularization is ineffective, architectural expansion is inefficient, orthogonality is too restrictive, and complex fusion is costly. In contrast, ELLA abandons the assumption that all interference must be uniformly suppressed and introduces the more effective principle of managing interference selectively. By regularizing only the high-magnitude, task-specific directions of the adapter space while permitting the reuse of low-magnitude, generalizable features, ELLA achieves a superior balance of stability and plasticity without the trade-offs that limit prior work, and integrates seamlessly into LoRA-based workflows, enabling scalable and robust CL adaptation for LLMs.
3 Method
3.1 Problem Formulation
Setup. In the supervised CL setting, a model encounters a stream of tasks $\{1,\dots,\mathcal{T}\}$ sequentially, where each task $t=\{(x^{t}_{i},y^{t}_{i})\}_{i=1}^{n_{t}}$ contains a labeled dataset $x^{t}_{i}\in\mathcal{X}_{t}$ and $y^{t}_{i}\in\mathcal{Y}_{t}$. Given a prediction model $h_{\Theta}$ with parameters $\Theta$, the objective is to maximize total log-likelihood over all tasks:
$\max_{\Theta}\sum_{k=1}^{\mathcal{T}}\sum_{(x,y)\in\mathcal{X}_{k},\mathcal{Y}_{k}}\log p_{\Theta}(y\mid x)\vskip-9.0pt$ (1)
Here, we consider a more challenging setting for ELLA: rehearsal-free CL with task-agnostic inference. (i) The model cannot store or access data from past tasks when learning a new one. (ii) At test time, the model must make predictions without knowing which task an input belongs to.
Figure 1: ELLA mitigates interference in continual LoRA training by accumulating past low-rank updates $\mathcal{W}_{\text{past}}$ and applying an energy-based alignment penalty $\Delta W_{t}*\mathcal{W}_{\text{past}}||_{F}^{2}$ to discourage overlap in high-magnitude, task-specific directions. This selective regularization enables parameter reuse in less-used subspaces, achieving a better trade-off between plasticity and stability without requiring task labels, data replay, or architectural modifications.
Low-Rank Adaptation (LoRA). Our method builds on LoRA Hu et al. (2022), a PEFT technique that leverages the observation that fine-tuning large pre-trained models (PTMs) can be effectively performed within a low-dimensional subspace Aghajanyan et al. (2020). Given a frozen pre-trained weight matrix $W_{\mathrm{init}}\in\mathbb{R}^{d\times k}$, LoRA introduces a low-rank update $\Delta W=AB$, where $A\in\mathbb{R}^{d\times r}$ and $B\in\mathbb{R}^{r\times k}$ with $r\ll\min(d,k)$. The original weights $W_{\mathrm{init}}$ remain unchanged during training, and only the low-rank matrices $A$ and $B$ are optimized. During the forward pass, the original operation $h=W_{\mathrm{init}}x$ becomes:
$h=W_{\mathrm{init}}x+\Delta Wx=W_{\mathrm{init}}x+ABx,\vskip-4.0pt$ (2)
This design reduces memory and compute overhead significantly, while achieving competitive performance to full fine-tuning across model families and benchmarks Houlsby et al. (2019).
3.2 Subspace-Aware Continual Adaptation
We propose ELLA, a simple yet effective extension to LoRA, as a CL framework that balances plasticity and stability in the adapter updates using a subspace-aware regularization strategy. Rather than enforcing strict orthogonality between LoRA updates, which can hinder forward transfer, ELLA penalizes overlap with past task-specific directions in proportion to their accumulated energy, thereby discouraging harmful interference while permitting reuse of low-magnitude directions that facilitate knowledge transfer, as illustrated in Fig. 1.
Let the LoRA update for task $t$ be $\Delta W_{t}=A_{t}B_{t}$, where $A_{t}\in\mathbb{R}^{d\times r}$ and $B_{t}\in\mathbb{R}^{r\times k}$. We construct a cumulative signal from the sum of LoRA-induced weight changes of past tasks:
$\mathcal{W}_{\text{past}}=\sum_{i=1}^{t-1}\Delta W_{i}.\vskip-8.0pt$ (3)
By construction, $\mathcal{W}_{\text{past}}$ encodes the dominant directions heavily used by prior tasks. Motivated by the observation that high-magnitude LoRA components are typically more task-discriminative Aghajanyan et al. (2020), we introduce the ELLA alignment penalty as follows:
$\mathcal{L}_{\text{ELLA}}=\|\Delta W_{t}*\mathcal{W}_{\text{past}}\|_{F}^{2},$ (4)
where $\|\cdot\|_{F}$ denotes the Frobenius norm. The overall training loss for task $t$ is then:
$\mathcal{L}=\sum_{(x,y)\in\mathcal{X}_{t},\mathcal{Y}_{t}}\log p_{\Theta}(y\mid x)+\lambda\cdot\mathcal{L}_{\text{ELLA}}.\vskip-5.0pt$ (5)
Here, $\lambda\geq 0$ controls the trade-off between plasticity (learning the new gradient step) and stability (suppressing interference with past subspaces).
Formal Characterization:
We next formalize the effect of this regularizer, as summarized in the following proposition. The detailed proof is discussed in Appendix A.
Proposition 1.
Let $G$ denote the unconstrained gradient step for task $t$, and let $E_{ij}=|(\mathcal{W}_{\text{past}})_{ij}|+\varepsilon$ denote the accumulated energy of past updates. Then the optimal update $\Delta W^{\star}_{t}$ under the ELLA-regularized objective is
$(\Delta W^{\star}_{t})_{ij}=\frac{G_{ij}}{1+\lambda E_{ij}^{2}}\vskip-9.0pt$ (6)
Moreover, the interference with past tasks is bounded by
$\bigl|\langle\Delta W^{\star}_{t},\mathcal{W}_{\text{past}}\rangle_{F}\bigr|\leq\frac{\|G\|_{F}}{2\sqrt{\lambda}}\;\|E^{-1}\odot\mathcal{W}_{\text{past}}\|_{F}\vskip-4.0pt$ (7)
At a high level, Eq. 6 shows that ELLA acts as an anisotropic shrinkage operator: coordinates with large past energy are shrunk more aggressively while low-energy coordinates remain flexible. The shrinkage strength is directly controlled by $\lambda$: larger values amplify the penalty on high-energy directions, prioritizing stability, while smaller values reduce the penalty, promoting plasticity. This property is further assessed in Sec. 5 and Fig. 6. The second result, Eq. 7, establishes a provable upper bound on task interference, ensuring that dominant past directions contribute minimally to forgetting. These properties formally justify ELLA’s ability to suppress destructive overlap while reusing underutilized subspaces for forward transfer.
4 Experiments
4.1 Datasets
Standard CL Benchmark (SC) is a popular CL benchmark designed for language models, comprising five text classification datasets introduced by Zhang et al. (2015). Following Wang et al. (2023a), we select AG News, Amazon Reviews, DBpedia, and Yahoo Answers, and form three shuffled task sequences, denoted as Orders $1$, 22, and $3$.
Long Sequence Benchmark (LS) extends the standard CL setup by incorporating $15$ tasks, including five classification datasets, nine GLUE and SuperGLUE benchmarks, and the IMDB dataset Razdaibiedina et al. (2023). Following the protocol of prior works Liao et al. (2025), we train each task using $1,000$ randomly selected samples, with $500$ samples per class reserved for testing. These tasks are also arranged into shuffled sequences Orders $4$, 55, and $6$.
TRACE is a benchmark tailored for CL in LLMs, covering a diverse set of $8$ tasks that span domains such as multiple-choice question answering, multilingual understanding, code generation, and mathematical reasoning Wang et al. (2023b). See App. B.2 for details on tasks and orderings.
4.2 Metrics
Let $a_{i,j}$ denote the testing performance on the $j$-th task after training on the $i$-th task. We evaluate across: Overall Accuracy (OA) Chaudhry et al. (2018): The average accuracy across all tasks after training on the last task, i.e., $\text{OA}_{\mathcal{T}}=\frac{1}{\mathcal{T}}\sum_{t=1}^{\mathcal{T}}a_{\mathcal{T},t}$; Forward Transfer (FWT) Lopez-Paz and Ranzato (2017): measures how much knowledge from previous tasks transfers to a new task, i.e., $\text{FWT}_{\mathcal{T}}=\frac{1}{\mathcal{T}}\sum_{t=1}^{\mathcal{T}}(a_{t,t}-a_{0,t})$, where $a_{0,t}$ refers to the performance of training task $t$ individually; Backward Transfer (BWT) Ke and Liu (2022): measures how much the learning of subsequent tasks influences the performance of prior tasks, i.e., $\text{BWT}_{\mathcal{T}}=\frac{1}{\mathcal{T}-1}\sum_{t=1}^{\mathcal{T}-1}(a_{\mathcal{T},t}-a_{t,t})$. Moreover, we also report general ability (GA) – the average generalization performance across unseen datasets post CL, and delta general ability (DeltaGA) – change in GA relative to the original LLM.
4.3 Baselines
We compare ELLA against a comprehensive suite of CL baselines. This includes naïve sequential fine-tuning approaches (SeqFT, SeqLoRA), classic regularization methods (EWC Kirkpatrick et al. (2017), LwF Li and Hoiem (2017)), and modern prompt-tuning techniques (L2P Wang et al. (2022b), LFPT5 Qin and Joty (2021)). We also benchmark against state-of-the-art adapter-based methods, including those that enforce strict orthogonality (O-LoRA Wang et al. (2023a), LB-CL Qiao and Mahdavi (2024)) as well as those that combine orthogonality with complex decomposition or replay mechanisms (DATA Liao et al. (2025), Recurrent KIF Feng et al. (2025)). Details of baselines are provided in Appendix B.3.
Methods Standard CL Benchmark (SC) Long Sequence Benchmark (LS) TRACE
Order 1 Order 2 Order 3 OA Order 4 Order 5 Order 6 OA Order 7 OA
T5-Large SeqFT de Masson D’Autume et al. (2019) 18.9 24.9 41.7 28.5 7.4 7.3 7.4 7.4 -
SeqLoRA 39.5 31.9 46.6 39.3 4.9 3.5 4.2 4.2 12.1
IncLoRA 63.4 62.2 65.1 63.6 63.0 57.9 60.4 60.5 -
SeqSVD 40.0 63.3 44.9 49.4 13.7 13.8 12.2 13.2 -
EWC Kirkpatrick et al. (2017) 46.3 45.3 52.1 47.9 44.9 44.0 45.4 44.8 -
LwF Li and Hoiem (2017) 52.7 52.9 48.4 51.3 49.7 42.8 46.9 46.5 -
L2P Wang et al. (2022b) 59.0 60.5 59.9 59.8 57.7 53.6 56.6 56.0 -
L-CL 75.3 73.5 71.9 73.6 66.5 64.0 69.0 66.5 -
B-CL 76.4 71.5 75.1 74.3 65.7 66.4 69.2 67.1 -
IncSVD 76.0 73.4 74.0 74.5 67.6 65.3 62.6 65.2 -
LB-CL Qiao and Mahdavi (2024) 76.9 76.5 76.8 76.7 68.4 67.3 71.8 69.2 -
O-LoRA Wang et al. (2023a) 73.5 71.4 70.0 71.6 65.4 65.2 65.2 65.3 23.1
   + MIGU Du et al. (2024) 77.1 77.0 75.6 76.6 67.3 68.5 74.0 70.0 -
DATA Liao et al. (2025) 71.5 70.5 68.0 70.0 71.5 70.5 68.0 70.0 16.7
 + Replay 77.0 75.6 75.2 75.9 75.6 73.2 74.1 74.3 36.5
LFPT5 Qin and Joty (2021) 66.6 71.2 76.2 71.3 69.8 67.2 69.2 68.7 -
SeqLoRAReplay 4.0 73.1 73.0 73.3 74.2 72.7 73.9 73.6 34.0
Recurrent-KIF Feng et al. (2025) - - - 78.4 - - - 77.8 -
ELLA (ours) 80.0 80.0 79.8 79.9 73.4 72.0 75.4 73.6 40.0
LLaMA3.1-8B SeqLoRA 75.88 74.40 74.35 74.86 67.81 65.93 63.80 65.85 28.23
O-LoRA Wang et al. (2023a) 69.39 67.58 71.44 69.46 69.54 64.42 66.50 66.82 28.45
DATA Liao et al. (2025) 76.10 75.69 75.52 75.77 71.34 70.77 72.95 71.69 31.33
   + Replay 77.56 77.01 76.03 76.83 73.10 73.05 74.17 73.44 34.16
SeqLoRAReplay 77.02 76.53 76.19 76.58 72.03 72.96 75.38 73.46 33.02
ELLA (ours) 77.80 77.20 77.70 77.57 72.87 72.84 76.82 74.18 33.29
Table 1: Overall Average Accuracy (OA) comparison of baselines and ELLA (ours) on Standard CL benchmark (Order 11, 22, 33) and Long Sequence benchmark (Order 44, 55, 66) and TRACE (Order 77) across multiple transfer orders. Methods in gray rely on replay mechanisms to boost performance. Best results in bold and second best underlined. $\star$ denotes methods that encounter OOM on GPUs smaller than $2\times$ H100100s, underscoring their limited scalability.
4.4 Implementation Details
We evaluate ELLA on a range of LLMs including encoder-decoder T55 models Raffel et al. (2020) (T55-Large, T55-XL), and decoder-only LLaMA models Touvron et al. (2023) (LLaMA-$3.1$ 88B). ELLA training is performed with the DeepSpeed library using V100100 16GB GPUs for T5 and A4040 48GB GPUs for LLaMA, while larger baselines such as DATA Liao et al. (2025) require H100100 80GB GPUs. All experiments are performed with instruction tuning Wei et al. (2021). Following Wang et al. (2023a), we apply LoRA modules (rank=$8$) to the Attention Q-V layers and report the average result of $3$ runs. More details in App. B.6.
4.5 Results
Table 1 reports overall performance across three CL benchmarks. Across all architectures, task orders, and evaluation settings, ELLA consistently sets a new state of the art in replay-free CL for LLMs. Remarkably, it not only dominates replay-free baselines but also surpasses strong replay-augmented methods, despite introducing no additional memory or computational burden.
On the Standard CL benchmark, ELLA achieves an average accuracy of $79.9$ on T55-Large, outperforming the strongest replay-free competitor Qiao and Mahdavi (2024) and exceeding DATA Liao et al. (2025) and Recurrent KIF Feng et al. (2025), both of which depend on data rehearsal. On the more demanding Long Sequence (LS) setup, ELLA improves over Qiao and Mahdavi (2024); Du et al. (2024) by upto $4.3\%$, demonstrating robustness under long horizons where interference typically compounds. This is further examined in Sec. 5 and Fig. 5. On TRACE, which stresses cross-domain CL adaptation, ELLA achieves a dramatic gain of up to +23.3+23.3, showing its ability to generalize to reasoning, code, and multilingual tasks.
Beyond average accuracy, Fig. 2 and 3 reveal deeper dynamics. ELLA achieves the best BWT, sharply reducing forgetting, while also obtaining the highest FWT, signaling effective knowledge reuse across tasks. This is crucial: unlike orthogonality-based approaches Wang et al. (2023a); Qiao and Mahdavi (2024), which guarantee stability but block positive transfer, ELLA explicitly allows low-energy subspace reuse, yielding consistent forward gains. Conversely, replay-based approaches achieve modest transfer but at the cost of large buffers and compute. Notably, ELLA sustains high performance on interference-sensitive tasks like QQP, IMDB, and DBpedia, where baselines degrade sharply. ELLA thus demonstrates that selective subspace decorrelation provides a more principled way to reconcile stability and plasticity than existing strategies.
Equally important, ELLA achieves these gains with no replay buffer, no generative augmentation, and no task identity signals. Replay-heavy methods Liao et al. (2025); Feng et al. (2025) not only incur large memory overheads but also encounter out-of-memory failures on LLaMA-3.1 8B when trained on GPUs smaller than H100100s. ELLA avoids these pitfalls entirely, remaining lightweight, scalable, and stable even at billion-parameter scale.
In summary, ELLA combines state-of-the-art replay-free accuracy with superior forward transfer, minimal forgetting, and low overhead, setting a new benchmark for scalable CL in LLMs.
4.6 Task Generalization
LLMs that are continually trained on new tasks often exhibit a decline in general performance, revealing catastrophic forgetting of their original capabilities. To assess this aspect, we evaluate the generalization ability of ELLA in the context of continual learning, focusing on its performance on five unseen cross-task benchmarks: MMLU Hendrycks et al. (2020), GSM88Cobbe et al. (2021), BBH Suzgun et al. (2022), AGIEval Zhong et al. (2023), and PIQA Bisk et al. (2020). More details on these datasets are provided in Appendix B.
We begin with a fine-tuned LLaMA 3.13.1-88B model trained on Order 1 of the SC Benchmark. As shown in Table 2, ELLA consistently achieves higher accuracy across all benchmarks compared to baselines, demonstrating its effectiveness in preserving knowledge while enabling strong cross-task generalization. Crucially, we see improved performance even over the original zero-shot model, underscoring its superiority in effectively combining acquired knowledge to address novel tasks. Beyond mitigating forgetting, ELLA delivers positive DeltaGA, showing that it not only retains prior skills but also facilitates forward transfer to unseen tasks. This indicates that ELLA provides a stable, rehearsal-free path to cross-task generalization, making it valuable for the continual deployment of LLMs in dynamic environments.
Figure 2: Performance impact on Order 44 (left) and 66 (right) in terms of BWT. We demonstrate superior resistance to performance decline than baselines (higher values indicate better retention of prior task performance).
Methods MMLU GSM8K BBH PIQA AGIEval DeltaGA
Zero Shot 59.64 69.67 35.65 72.42 43.92 0
SeqLoRA 56.05 70.20 35.16 67.85 41.02 -11.02
O-LoRA 57.36 66.79 33.82 64.31 44.21 -14.81
SeqLoRAReplay 55.81 66.43 32.90 67.77 41.58 -16.81
DATA 46.84 14.17 12.02 59.02 39.11 -110.14
Ours 59.92 71.05 36.37 75.52 43.92 5.48
Table 2: Generalization ability (GA) on unseen tasks.
Figure 3: Performance comparison across different backbone size and model families.
4.7 Scale to Larger Models
To assess the scalability and robustness of ELLA across model sizes and architectures, we examine CL performance across T55-Base (220220M), T55-Large (770770M), T55-XL (33B), and LLaMA-$3.1$ 88B on Order $5$ of the Long Sequence Benchmark. As seen in Fig. 3, ELLA demonstrates consistent gains in average performance and transfer metrics as model size increases, outperforming baselines across all backbone configurations.
It is worth noting that even with the largest backbone model, O-LoRA(88B) still falls short in terms of FWT compared to the smallest version of ELLA (220220M). This further highlights the crucial importance of selecting the pertinent PEFT algorithm for continual adaptation, rather than relying solely on scaling backbone size. Notably, the T55-XL model achieves higher OA than the larger LLaMA-$3.1$ model, highlighting that encoder-decoder architectures like T55 might be more effective for CL on classification tasks than decoder-only LLMs. We attribute this to the full-sequence bidirectional attention and cross-attention present in encoder-decoder designs, which better facilitate knowledge retention and task transfer than the causally masked architecture of models like LLaMA. These findings underscore that, beyond scaling size, architectural choice plays a crucial role, and shows that forgetting is model-dependent.
Method Trainable Params Storage (MB) Replay Time/Epoch (mins)
SeqLoRA 0.062 0 0 4
O-LoRA 0.062 31.46 0 4.5
SeqLoRAReplay 0.062 0 2% 4
DATA 0.369 147.46 2% 6.5
Ours 0.062 4.19 0 4.5
Table 3: Comparison of training efficiency, memory overhead, and storage requirements with T55-Large.
Figure 4: Histogram of prediction loss changes after training on a new task. The ELLA constraint helps reduce the changes – preserve the loss of previous tasks – in comparison to when it is not present ($\lambda=0$).
4.8 Efficiency Analysis
As shown in Table 3, ELLA matches O-LoRA in trainable parameters while cutting storage from 31.4631.46MB to just 4.194.19MB–an $8\times$ reduction. Unlike replay-based methods such as DATA and SeqLoRAReplay, ELLA requires no buffer or feature storage and adds only minimal runtime cost. By embedding a lightweight regularization penalty into the loss instead of storing multiple task-specific modules, ELLA achieves strong efficiency in memory, compute, and training time – showing that ELLA is scalable and practical for long-horizon CL, where resource demands are typically prohibitive.
5 Discussions
Does ELLA preserve previous task performance during CL? We examine the effectiveness of ELLA in preventing degradation on previously learned tasks by measuring the change in prediction loss on past-task batches after training a new task. As shown in Fig. 4, ELLA significantly reduces the number of batches experiencing large increases in loss, especially in the high-loss tail region. This indicates that ELLA better preserves useful gradients from prior tasks by leveraging its alignment penalty strategy. In contrast, the baseline (without ELLA) exhibits a broader distribution of loss spikes, revealing higher susceptibility to catastrophic forgetting. These results confirm that ELLA enhances stability across tasks without requiring replay or task labels.
Directional Consistency of Updates Over Task Sequence. In Fig. 5, we analyze the degree of opposing weight updates after each new task for Order $5$ on both T55-Large and LLaMA3.13.1-88B. Specifically, we measure the magnitude of weight change that occurs in the direction opposite to that of the previous task, indicating a misalignment between updates over time. Standard LoRA exhibits consistently high opposing updates, suggesting that new learning often disrupts previously acquired representations, while ELLA significantly reduces such opposing direction weight updates, enabling smoother and more stable knowledge accumulation across tasks. This improvement is consistent across both encoder-decoder and decoder-only models.
Figure 5: Opposing direction weight change across task sequence for T55-Large (left) and LLaMA-$3.1$ 88B (right). ELLA consistently reduces backward-conflicting updates, promoting stable continual adaptation.
Impact of $\lambda$ Scaling. Fig. 6 shows that $\lambda=0$ (i.e, a vanilla LoRA) results in severe forgetting, as evidenced by low OA and BWT, while increasing $\lambda$ progressively improves both metrics by constraining interference. Excessively large values, however, hinder new task adaptation, lowering OA. The best performance arises at moderate $\lambda$, where the two terms in Eq. 5 are balanced, yielding both stability and plasticity. This behavior is consistent with our theoretical characterization in Appendix A, which shows that $\lambda$ acts as an anisotropic shrinkage factor in our update, interpolating between unconstrained learning and stability-preserving adaptation.
Figure 6: Impact of $\lambda$ scaling on Order 11. Overall Accuracy and BWT as $\lambda$ varies shows that moderate scaling achieves the best stability-plasticity balance.
LoRA_dim Order 1 Order 2 Order 3 Avg
2 72.29 74.00 77.08 74.46
4 73.22 75.15 77.72 75.36
8 79.95 80.00 79.82 79.92
16 77.38 77.65 76.19 77.07
Table 4: Impact of LoRA rank on SC Benchmark.
Studying Optimal LoRA Rank for Plasticity-Stability Tradeoff. We investigate how LoRA rank influences CL by evaluating ELLA with T55-Large across three task orders while varying the LoRA dimension $r$. As shown in Table 4, performance improves with increasing $r$, peaking at $r=8$. Very low ranks ($r=2$) severely limit plasticity and hinder adaptation to new tasks, while higher ranks ($r=16$) reduce stability by overfitting to current tasks. This underscores that careful rank selection is crucial for sustaining continual adaptation without catastrophic forgetting.
6 Conclusion
We introduced ELLA, a simple yet powerful approach for continual adaptation of LLMs that mitigates forgetting without replay or task identifiers. By encouraging controlled de-alignment from past subspaces, ELLA reduces destructive interference while enabling reuse of underutilized directions. We provide formal guarantees showing that ELLA’s update admits a closed-form solution with bounded interference, offering a principled foundation for balancing plasticity and stability. Extensive experiments across benchmarks, model architectures, and scales demonstrate that ELLA consistently improves both forward and backward transfer, while remaining highly parameter- and memory-efficient. Notably, ELLA surpasses state-of-the-art baselines and scales robustly to larger backbones such as T5-33B and LLaMA-3.13.1-88B with accuracy gains of up to $9.6\%$. It also uniquely enhances generalization performance on previously unseen tasks compared to the original LLM models – an ability absent in prior methods. These findings position ELLA as a lightweight, scalable, and universal method for lifelong adaptation in LLMs, demonstrating that efficiency and continual improvement need not come at the cost of forgetting.
7 Limitations
While our approach demonstrates strong performance, its scalability to more complex continual learning scenarios involving hundreds of tasks remains an open question. Moreover, although task identification is not required at inference time, our current training still assumes task labels to assign task-specific LoRA parameters. Developing task-agnostic training strategies is a promising direction for future work. Finally, due to resource constraints, we have not evaluated our method on larger models like LLaMA-3.13.1-7070B, and extending ELLA to continual multimodal LLMs remains an exciting avenue for exploration.
8 Ethical Considerations
ELLA represents a significant step towards practical and scalable lifelong learning for LLMs. By enabling efficient, replay-free updates, our work makes it more feasible to keep models current in dynamic environments without hindering data privacy and reducing the environmental footprint of AI systems. While ELLA’s design promotes efficiency and privacy, it does not inherently mitigate the risks of misuse, such as the generation of biased or harmful content, which are endemic to the base models. Deploying models updated with ELLA requires a continued commitment to responsible data curation, fairness audits, and robust safety mechanisms to address these challenges. The datasets and models we used are public, hence there are no privacy issues.
9 AI Writing Statement
This paper utilized AI assistance for language polishing of the manuscript, including vocabulary correction and spell checking.
Appendix
This appendix provides supplementary material to support the findings presented in the main paper. We begin in Section A with a detailed theoretical analysis of the ELLA regularizer, offering a formal proof of its properties, including its formulation as an anisotropic shrinkage operator and its ability to bound interference provably. In Section B, we provide a comprehensive overview of our experimental settings. This includes detailed descriptions of the datasets, baselines, task sequences, instruction prompts, and specific hyperparameter configurations used in our evaluations. Finally, Section C lists and discusses the artifact licenses.
Appendix A Theoretical Analysis of the ELLA Regularizer
Our theoretical analysis shows that the ELLA regularizer has two key properties: (i) its optimal update is an anisotropic shrinkage operator that selectively preserves plasticity, and (ii) it establishes a provable bound on the interference with past tasks, guaranteeing stability.
A.1 Setup and Preliminaries
Our analysis is set in the context of continual learning, where a model adapts to a sequence of tasks. For each new task $t$, the model is adapted by learning a low-rank update matrix, $\Delta W_{t}$, using the LoRA method. This update is defined as the product of two smaller matrices:
$\Delta W_{t}=A_{t}B_{t}$ (8)
where $A_{t}\in\mathbb{R}^{d\times r}\text{ and }B_{t}\in\mathbb{R}^{r\times k}\text{ with rank }r\ll\min\{d,k\}$. This matrix represents the specific knowledge acquired for task $t$. We define the aggregated matrix of past updates as:
$\mathcal{W}_{past}=\sum_{i=1}^{t-1}\Delta W_{i}$ (9)
The practical regularization term used in our work is $\|\mathcal{W}_{past}\odot\Delta W_{t}\|_{F}^{2}$. For analytical tractability, we replace this term with a closely related objective. We introduce an energy matrix $E\in\mathbb{R}^{d\times k}$ derived from past updates, be defined as $E_{ij}=|(\mathcal{W}_{past})_{ij}|+\varepsilon$ for a small constant $\varepsilon>0$ that ensures all entries are positive and $E^{-1}$ is well-defined. This preserves the core principle of suppressing alignment with high-energy coordinates.
For the current task $t$, there is an ideal, unconstrained update that would best minimize the task loss, $\mathcal{L}_{t}$. This is the standard gradient step, which we denote as $G$.
$G\triangleq\nabla\mathcal{L}_{t}(W_{t-1})$
We think of $G$ as the update we would make if we did not have to worry about forgetting past knowledge.
The goal of ELLA is to find an optimal update $\Delta W_{t}$ that is close to the target gradient $G$ (learning the new task) but is penalized for changing parameters that were important for past tasks (not forgetting). This is formulated as the following regularized optimization problem:
$\min_{\Delta W}\;\;\underbrace{\frac{1}{2}\|\Delta W_{t}-G\|_{F}^{2}}_{\text{Plasticity Term}}\;+\;\underbrace{\frac{\lambda}{2}\,\|E\odot\Delta W_{t}\|_{F}^{2}}_{\text{Stability Term}}$ (10)
where $\lambda>0$ is the regularization strength and $\odot$ denotes the element-wise Hadamard product. The plasticity term encourages $\Delta W$ to be similar to $G$, while the stability term penalizes changes to coordinates with high energy (large values in $E$).
Proposition 1 The solution $\Delta W_{t}^{\star}$ to the ELLA objective in Eq. (10) has the following properties:
  1. (i)
    It is an anisotropic shrinkage operator applied to the unconstrained step $G$, with the closed-form solution:
    $(\Delta W_{t}^{\star})_{ij}=\frac{G_{ij}}{\,1+\lambda E_{ij}^{2}\,}$
  2. (ii)
    The interference with past updates, measured by the inner product $\langle\Delta W_{t}^{\star},\mathcal{W}_{past}\rangle_{F}$, is bounded as follows:
    $\bigl|\langle\Delta W_{t}^{\star},\mathcal{W}_{past}\rangle_{F}\bigr|\leq\frac{\|G\|_{F}}{2\sqrt{\lambda}}\;\|E^{-1}\odot\mathcal{W}_{past}\|_{F}$
Proof.
Proof of part (i): The objective function in Eq. (10) is strongly convex since its second derivative is strictly positive, and separable across the matrix entries $(i,j)$. We can therefore solve for each entry $(\Delta W_{t})_{ij}$ independently. Let $z=(\Delta W_{t})_{ij}$, $g=G_{ij}$, and $e=E_{ij}$. The per-coordinate objective is:
$\min_{z\in\mathbb{R}}\;\;\tfrac{1}{2}(z-g)^{2}+\tfrac{\lambda}{2}e^{2}z^{2}$
Taking the derivative with respect to $z$ and setting it to zero yields the optimal solution $z^{\star}$:
$(z^{\star}-g)+\lambda e^{2}z^{\star}=0\quad\Rightarrow\quad(1+\lambda e^{2})\,z^{\star}=g$
This gives the closed-form solution for each coordinate, proving part (i):
$(\Delta W_{t}^{\star})_{ij}=\frac{G_{ij}}{\,1+\lambda E_{ij}^{2}\,}$ (11)
This operator selectively shrinks the update for each coordinate based on the accumulated energy of past updates $E_{ij}$, with larger energy leading to stronger shrinkage.
Proof of part (ii): Next, we first bound the norm of the energy-weighted update. Using the solution from Eq. (11):
$\|E\odot\Delta W_{t}^{\star}\|_{F}^{2}$ $=\sum_{ij}\frac{E_{ij}^{2}\,G_{ij}^{2}}{(1+\lambda E_{ij}^{2})^{2}}$ (12)
$\leq\left(\sup_{x\geq 0}\frac{x}{(1+\lambda x)^{2}}\right)\,\sum_{ij}G_{ij}^{2}$ (13)
The supremum is attained at $x=1/\lambda$, yielding $\sup_{x\geq 0}\frac{x}{(1+\lambda x)^{2}}=\frac{1}{4\lambda}$. Therefore,
$\|E\odot\Delta W_{t}^{\star}\|_{F}^{2}\leq\frac{1}{4\lambda}\,\|G\|_{F}^{2}$ (14)
This shows that the effective penalty is never worse than a $\tfrac{1}{4\lambda}$ fraction of the gradient energy, ensuring the shrinkage remains controlled.
Now, we try to bound the interference term $\langle\Delta W_{t}^{\star},\mathcal{W}_{past}\rangle_{F}$. By rewriting the inner product and applying the Cauchy-Schwarz inequality:
$\bigl|\langle\Delta W_{t}^{\star},\mathcal{W}_{past}\rangle_{F}\bigr|$ (15)
$=\bigl|\langle E\odot\Delta W_{t}^{\star},\,E^{-1}\odot\mathcal{W}_{past}\rangle_{F}\bigr|$ (16)
$\leq\|E\odot\Delta W_{t}^{\star}\|_{F}\ \|E^{-1}\odot\mathcal{W}_{past}\|_{F}$ (17)
Substituting the bound from Eq. (14) into Eq. (17) yields the final interference bound:
$\bigl|\langle\Delta W_{t}^{\star},\mathcal{W}_{past}\rangle_{F}\bigr|\leq\frac{\|G\|_{F}}{2\sqrt{\lambda}}\;\|E^{-1}\odot\mathcal{W}_{past}\|_{F}$
This completes the proof of part (ii). The term $\|E^{-1}\odot\mathcal{W}_{past}\|_{F}$ ensures that coordinates with high energy (large $E_{ij}$) contribute minimally to the bound, formally proving that ELLA enforces stability by suppressing interference in high-energy coordinates, while preserving plasticity where capacity remains.
Appendix B Experimental Settings
B.1 Datasets
Train Tasks. Tables 5 and  6 provide detailed information on the datasets utilized in our continual learning (CL) experiments. Table 5 presents the $15$ datasets included in the Long Sequence Benchmark Razdaibiedina et al. (2023), while Table 6 outlines the $8$ datasets from TRACE Wang et al. (2023b). Both tables include the corresponding evaluation metrics for each dataset.
Unseen tasks. We select the following datasets to test generalization performance post continual adaptation: (1) Multitask Language Understanding (MMLU) Hendrycks et al. (2020), which includes multiple-choice questions across $57$ subjects. (2) GSM8K Cobbe et al. (2021), which is a high-quality linguistically diverse multi-step elementary math reasoning dataset. (3) BIG-Bench Hard (BBH) Suzgun et al. (2022), which includes $27$ challenging tasks spanning arithmetic, symbolic reasoning, and more, derived from BIG-Bench (BB) Srivastava et al. (2022). Most of the data consists of multiple-choice questions. (4) AGIEval Zhong et al. (2023), which includes a wide range of high-quality official entrance exams, qualifying exams, and advanced competitions tailored to human participants. (5) PIQA Bisk et al. (2020) which is a dataset for commonsense reasoning, and was created to investigate the physical knowledge of existing models in NLP.
Dataset Name Category Task Domain Metric
Yelp CL Benchmark Sentiment Analysis Yelp Reviews Accuracy
Amazon CL Benchmark Sentiment Analysis Amazon Reviews Accuracy
DBPedia CL Benchmark Topic Classification Wikipedia Accuracy
Yahoo CL Benchmark Topic Classification Yahoo Q&A Accuracy
AG News CL Benchmark Topic Classification News Accuracy
MNLI GLUE Natural Language Inference Various Accuracy
QQP GLUE Paragraph Detection Quora Accuracy
RTE GLUE Natural Language Inference News, Wikipedia Accuracy
SST-2 GLUE Sentiment Analysis Movie Reviews Accuracy
WiC SuperGLUE Word Sense Disambiguation Lexical Databases Accuracy
CB SuperGLUE Natural Language Inference Various Accuracy
COPA SuperGLUE Question and Answering Blogs, Encyclopedia Accuracy
BoolQA SuperGLUE Boolean Question and Answering Wikipedia Accuracy
MultiRC SuperGLUE Question and Answering Various Accuracy
IMDB SuperGLUE Sentiment Analysis Movie Reviews Accuracy
Table 5: The details of 15 classification datasets in the Long Sequence Benchmark Razdaibiedina et al. (2023). The first five tasks correspond to the standard CL benchmark Zhang et al. (2015).
Dataset Source Avg len Metric Language #Data
Domain-specific
ScienceQA Science 210 Accuracy English 5,000
FOMC Finance 51 Accuracy English 5,000
MeetingBank Meeting 2853 ROUGE-L English 5,000
Multi-lingual
C-STANCE Social media 127 Accuracy Chinese 5,000
20Minuten News 382 SARI German 5,000
Code Completion
Py150 Github 422 Edim Similarity Python 5,000
Mathematical Reasoning
NumGLUE-cm Math 32 Accuracy English 5,000
NumGLUE-ds Math 21 Accuracy English 5,000
Table 6: The overview of dataset statistics in TRACE Wang et al. (2023b). The ‘Source’ indicates the origin of the context. ‘Avg len’ denotes the average length, measured in word count for English, German, and code datasets, and in character count for Chinese datasets. ‘SARI’ is a score that is specific to evaluating simplification tasks.
Benchmark Order Task Sequence
Standard CL Benchmark 1 dbpedia $\rightarrow$ amazon $\rightarrow$ yahoo $\rightarrow$ ag
2 dbpedia $\rightarrow$ amazon $\rightarrow$ ag $\rightarrow$ yahoo
3 yahoo $\rightarrow$ amazon $\rightarrow$ ag $\rightarrow$ dbpedia
Long Sequence Benchmark 4 mnli $\rightarrow$ cb $\rightarrow$ wic $\rightarrow$ copa $\rightarrow$ qqp $\rightarrow$ boolqa $\rightarrow$ rte $\rightarrow$ imdb $\rightarrow$ yelp $\rightarrow$ amazon $\rightarrow$ sst-2 $\rightarrow$ dbpedia $\rightarrow$ ag $\rightarrow$ multirc $\rightarrow$ yahoo
5 multirc $\rightarrow$ boolqa $\rightarrow$ wic $\rightarrow$ mnli $\rightarrow$ cb $\rightarrow$ copa $\rightarrow$ qqp $\rightarrow$ rte $\rightarrow$ imdb $\rightarrow$ sst-2 $\rightarrow$ dbpedia $\rightarrow$ ag $\rightarrow$ yelp $\rightarrow$ amazon $\rightarrow$ yahoo
6 yelp $\rightarrow$ amazon $\rightarrow$ mnli $\rightarrow$ cb $\rightarrow$ copa $\rightarrow$ qqp $\rightarrow$ rte $\rightarrow$ imdb $\rightarrow$ sst-2 $\rightarrow$ dbpedia $\rightarrow$ ag $\rightarrow$ yahoo $\rightarrow$ multirc $\rightarrow$ boolqa $\rightarrow$ wic
TRACE 7 c-stance $\rightarrow$ fomc $\rightarrow$ meetingbank $\rightarrow$ py150 $\rightarrow$ scienceqa $\rightarrow$ numglue-cm $\rightarrow$ numglue-ds $\rightarrow$ 20minuten
Table 7: Seven distinct orders of task sequences were employed for the experiments in continual learning. Orders 1-3 align with the Standard CL Benchmarks, as adopted in previous studies Liao et al. (2025). Orders 4-6 pertain to the Long Sequence Benchmarks, which encompass a total of 15 tasks Razdaibiedina et al. (2023). Order 7 refers to the TRACE benchmark, specifically designed for LLMs, and comprises eight datasets Wang et al. (2023b).
B.2 Task Sequence Orders
We report all task orders used for our CL experiments in Table 7.
B.3 Baselines
We compare our method against a comprehensive set of recent CL baselines, detailed as follows: (1) SeqFT sequentially fine-tunes all model parameters without CL mechanisms like regularization or replay (RorR). (2) SeqLoRA applies fixed-size LoRA tuning per task with/without replay. (3) IncLoRA allows incremental learning of new LoRA parameters for each task without constraints. (4) SeqSVD uses fixed-rank SVD adapters trained sequentially without RorR. (5) EWC Kirkpatrick et al. (2017) finetunes LoRA with a regularization loss to prevent interference with past tasks. (6) LwF Li and Hoiem (2017) distills the model of the last task using the current task data. (7) L2P Wang et al. (2022b) instantiates a prompt pool for adaptive prompt selection and prompt tuning for individual samples. (8) LFPT5 Qin and Joty (2021) learns soft prompts and generates pseudo-data for replay. (9) L-CL trains SVD adapters incrementally with SVD regularization. (10) B-CL applies SVD regularization with gradient projection. (11) O-LoRA Wang et al. (2023a) enforces orthogonality between LoRA updates across tasks. (12)  MIGU Du et al. (2024) updates parameters based on gradient magnitude. (13) LB-CL Qiao and Mahdavi (2024) initializes low-rank matrix parameters in new tasks from specific past parameters besides enforcing orthogonality. (14) DATA Liao et al. (2025) decomposes attention into high and low rank subspaces, and leverages orthogonality with optional replay. (15) Recurrent KIF Feng et al. (2025) uses an expensive recurrent mechanism to modulate task-specific adapter reuse using replay.
Task Prompts
NLI What is the logical relationship between the "sentence 1" and the "sentence 2"?
Choose one from the option.
QQP Whether the "first sentence" and the "second sentence" have the same meaning?
Choose one from the option.
SC What is the sentiment of the following paragraph? Choose one from the option.
TC What is the topic of the following paragraph? Choose one from the option.
BoolQA According to the following passage, is the question true or false?
Choose one from the option.
MultiRC According to the following passage and question, is the candidate answer true
or false? Choose one from the option.
WiC Given a word and two sentences, whether the word is used with the same sense
in both sentence? Choose one from the option.
FOMC What is the monetary policy stance for the following text?
Choose one from the option.
20Minuten Provide a simplified version of the following paragraph in German.
ScienceQA Choose an answer for the following question and give your reasons.
NumGLUE-cm Solve the following math problem.
NumGLUE-ds Solve the following math problem.
Py150 Continue writing the code.
MeetingBank Write a summary of the following meeting transcripts.
C-STANCE Determine the attitude of the following text towards the specified object.
Choose one from the option.
Table 8: Instructions for different tasks.
B.4 Instruction Tuning.
Instruction-following is a fundamental capability for LLMs to serve as an effective interface between humans and AI systems Wei et al. (2021); Ouyang et al. (2022). We adopt instruction tuning as our training paradigm for two key reasons: (1) it enables the incorporation of human prior knowledge via explicit instructions, thereby facilitating more efficient learning; and (2) it enhances generalization by guiding the model to learn underlying principles that apply across tasks.
All tasks are formatted using a unified schema consisting as follows: (1) Task Instruction – a clear description of how to map an input (e.g., a sentence or document) to an output; (2) Options – the constrained set of valid output labels for the task; (3) Text – the input instance provided to the model; and (4) Answer – the correct target output. This combined sequence is fed into the pretrained model to guide prediction for all experiments.
B.5 Task Instructions
Table 8 shows the prompts used for different tasks. NLI denotes natural language inference, and includes tasks MNLI, RTE and CB. SC denotes sentiment analysis, including Amazon, Yelp, SST-2 and IMDB. TC denotes topic classification and contains the tasks AG News, Dbpedia and Yahoo.
B.6 Implementation Details
Our implementation for ELLA uses PyTorch v2.0.1 Paszke et al. (2019) and Deepspeed v0.10.0 Rasley et al. (2020). Our experiments were conducted on machines equipped with 44 A4040 GPUs/ 88 V100100 GPUs for ELLA and all baselines, except DATA/DATA+Replay, which required 22 H100100 GPUs. For all orders of task streams for the Standard-CL Benchmark and the Long Sequence Benchmark, we trained the models with one epoch using the AdamW optimizer Loshchilov and Hutter (2017) and WarmupLR scheduler 10 with a total batch size of $32$. We used a constant learning rate of 1e31e-3 for T55 and 1e41e-4 for LLaMA, a dropout rate of $0.1$, and a weight decay rate of 0. For TRACE Order $7$ (C-STANCE, FOMC, MeetingBank, Py150, ScienceQA, NumGLUE-cm, NumGLUE-ds, 20Minuten), we trained with $5000$ samples $5$, 33, 77, 55, 33, 55, 55, 77 epochs respectively. The coefficient $\lambda$ for different task orders has been reported in Table 9. $\lambda$ was selected based on performance on a small, held-out validation set. To establish an effective search range, we followed the common heuristic of choosing values that scale the regularization term, $\mathcal{L}_{\text{ELLA}}$ (promoting stability), to a similar order of magnitude as the current task accuracy loss (promoting plasticity) for balanced learning. We observed that performance was robust across this range and that a single $\lambda$ often generalized well across multiple subsequent tasks. The LoRA rank was set to $8$ for all experiments, and the proportion of past task data mixed in for replay methods was set to $2\%$ of the original training set.
Order T55 $\lambda$ LLaMA $\lambda$
1–3 $0,3\times 10^{4},\,\ldots,\,3\times 10^{4}$ $0,3\times 10^{6},\,\ldots,\,3\times 10^{6}$
[…]
Table 9: Coefficient $\lambda$ settings for different task orders in T55 and LLaMA.
Appendix C Artifact Licenses
According to their license, all the LLMs used in this paper fall under acceptable use cases. The licenses are listed for perusal: T55-Base (https://huggingface.co/google-t5/t5-base), T55-Large (https://huggingface.co/google-t5/t5-large), T55-XL (https://huggingface.co/google/t5-v1_1-xl), LLaMA-3.1-8b (https://huggingface.co/meta-llama/Llama-3.1-8B/blob/main/LICENSE).